# Poison-splat: Computation Cost Attack on 3D Gaussian Splatting

### v1

@@ Abstract
3D Gaussian splatting (3DGS), known for its groundbreaking performance and efficiency, has become a dominant 3D representation and brought progress to many 3D vision tasks. However, in this work, we reveal a significant security vulnerability that has been largely overlooked in 3DGS: the computation cost of training 3DGS could be maliciously tampered by poisoning the input data. By developing an attack named Poison-splat, we reveal a novel attack surface where the adversary can poison the input images to drastically increase the computation memory and time needed for 3DGS training, pushing the algorithm towards its worst computation complexity. In extreme cases, the attack can even consume all allocable memory, leading to a Denial-of-Service (DoS) that disrupts servers, resulting in practical damages to real-world 3DGS service vendors. Such a computation cost attack is achieved by addressing a bi-level optimization problem through three tailored strategies: attack objective approximation, proxy model rendering, and optional constrained optimization. These strategies not only ensure the effectiveness of our attack but also make it difficult to defend with simple defensive measures. We hope the revelation of this novel attack surface can spark attention to this crucial yet overlooked vulnerability of 3DGS systems [^1] Our code is available at https://github.com/jiahaolu97/poison-splat..
@@ 1 Introduction
We develop a computation cost attack method named Poison-splat as a proof-of-concept of this novel attack vector. Poison-splat takes the form of training data poisoning (Tian et al., 2022), where the attacker manipulates the input data fed into the victim 3DGS system. This is practical in real-world scenarios, as commercial 3D service providers like Kiri (KIRI, ), Polycam (Polycam, ) and Spline (Spline, ) receive images or videos from end users to generate 3D captures. An attacker can hide themselves among normal users to submit poisoned data and launch the attack stealthily, or even secretly manipulate data uploaded by other users. During peak usage periods, this attack competes for computing resources with legitimate users, slowing down service response time and potentially causing severe consequences like service crashes, leading to financial losses for service providers.
@@ 4 Experiments
Implementation details. In our experiments, we use the official implementation of 3D Gaussian Splatting [^5] https://github.com/graphdeco-inria/gaussian-splatting (Kerbl et al., 2023) to implement the victim behavior. Following their original implementation, we use the recommended default hyper-parameters, which were proved effective across a broad spectrum of scenes from various datasets. For testing black-box attack effectiveness, we use the official implementation of Scaffold-GS [^6] https://github.com/city-super/Scaffold-GS (Lu et al., 2024) as a black box victim. All experiments reported in the paper were carried out on a single NVIDIA A800-SXM4-80G GPU.
@@ 4.1 Main Results
We report the comparisons of the number of Gaussians, peak GPU memory, training time and rendering speed between clean and poisoned data across three different datasets in Tables 1 to show the effectiveness of our attack. The full results containing all scenes are listed in the Appendix Table 3 and Table 4. We provide visualizations of example poisoned datasets and corresponding victim reconstructions in Appendix Section E.
@@ 4.2 Black-box Attack Performance
The answer is affirmative. To test the effectiveness of our attack on a black-box victim, we chose Scaffold-GS (Lu et al., 2024), a 3DGS variant algorithm designed with a focus on efficiency. Scaffold-GS differs significantly from traditional Gaussian Splatting as it utilizes anchor points to distribute local 3D Gaussians, which greatly alters the underlying feature representation. We directly tested Scaffold-GS on the poisoned dataset we crafted for the traditional 3DGS algorithm, making the victim (Scaffold-GS trainer) a black-box model. The results, as presented in Table 2 and Table 5, demonstrate the attack’s capability to generalize effectively, as evidenced by increased Gaussian parameters, higher GPU memory consumption, and reduced training speed.
@@ 5 Conclusion and Discussions
Societal Impacts. While our methods could potentially be misused by malicious actors to disrupt 3DGS service providers and cause financial harm, our aim is not to facilitate such actions. Instead, our objective is to highlight significant security vulnerabilities within 3DGS systems and prompt the community—researchers, practitioners, and providers—to recognize and address these gaps. We aim to inspire the development of more robust algorithms and defensive strategies, enhancing the security and reliability of 3DGS systems in practical applications. We are committed to ethical research and do not endorse using our findings to inflict societal harm.
@@ 6 Acknowledgement
The Appendix is organized as follows: Section A discusses related works; Section B provides a formal and comprehensive description of 3D Gaussian Splatting training. Section C offers the full results of experiments mentioned in the main paper, presented here due to page constraints. Section D lists all the notations used throughout the paper. Section E presents examples of attacker-crafted poisoned images and the corresponding reconstructions from victim models trained with those images.
@@ Appendix A Related Works
3D Gaussian Splatting is receiving overwhelming popularity in the 3D vision domain (Chen & Wang, 2024; Fei et al., 2024; Wu et al., 2024b; Dalal et al., 2024). The explicit nature of 3D Gaussian Splatting enables real-time rendering capabilities and unparalleled levels of control and editability, demonstrating its effectiveness in 3D generation. Since its introduction in (Kerbl et al., 2023), 3D Gaussian Splatting has facilitated a diverse array of downstream applications, including text-to-3D generation (Chen et al., 2024; Tang et al., 2024; Yi et al., 2024), Simultaneous Localization and Mapping (Keetha et al., 2024; Matsuki et al., 2024; Yan et al., 2024), human avatars (Kocabas et al., 2024; Lei et al., 2024; Zielonka et al., 2023), 4D-dynamic scene modeling (Luiten et al., 2024; Wu et al., 2024a; Yang et al., 2024), and other applications (Xie et al., 2024b; Shen et al., 2024; Ye et al., 2023; Wang et al., 2024b; a).
@@ A.2 Attacks on Machine Learning Systems
Machine learning has been shown to be vulnerable to various types of attacks (Pitropakis et al., 2019; Rigaki & Garcia, 2023; Chakraborty et al., 2018; Oliynyk et al., 2023; Tian et al., 2022). The majority of attacks target the either confidentiality (including membership inference (Shokri et al., 2017; Salem et al., 2019; Choquette-Choo et al., 2021), data reconstruction attacks (Fredrikson et al., 2015; Carlini et al., 2021; 2023; Yu et al., 2023; Geiping et al., 2020; Lu et al., 2022) and model stealing attacks (Tramèr et al., 2016; Chandrasekaran et al., 2020)) or integrity (like adversarial attacks (Biggio et al., 2013; Szegedy et al., 2013; Dong et al., 2018; 2019) and data poisoning attacks (Barreno et al., 2010; Jagielski et al., 2018; Biggio et al., 2012; Mei & Zhu, 2015b)). Different from the above mentioned attacks, we aim to illustrate that by carefully crafting input images, the attacker can indeed attack availability, i.e., timely and cost-affordable access to machine learning service. The most similar research is the work by Shumailov et al. (2021), which is the first Denial-of-Service (DoS) attack towards machine learning systems. It proposed sponge examples which can cause much more latency and energy consumption for language models and image classifiers, driving deep learning models to their worst performance. Different from the work (Shumailov et al., 2021), our attack targets 3D reconstruction algorithms and is not driven by neural networks.
A.2.1 Data Poisoning Attack
@@
We consider our proposed attacks to fall into the category of data poisoning attack, which manipulates training data to launch the attack at the training stage. The attacker makes the first move to craft poisoned training samples, and the victim has a goal of training the system based on samples provided by the attacker. Solving this problem typically involves solving a max-min bi-level optimization problem, where the attacker wants to maximize some adversarial loss (i.e., test error or computational cost) and victim wants to minimize the training objective (i.e., training loss). Different literature solves the problem using different techniques (Biggio et al., 2012; Mei & Zhu, 2015b; a; Li et al., 2016). For many machine learning systems, given the inherent complexity and their reliance on large datasets, it is intractable to exactly do the bi-level optimization. Muñoz-González et al. (2017) exploited back-gradient optimization technique to approximate real gradients through reverse-mode differentiation.  Huang et al. (2020) formulated as a meta-learning problem, and approximated the inner optimal solution by training only $K$ SGD steps for each outer objective evaluation. Koh et al. proposed three methods to approximate the expensive bi-level optimization: influence functions (Koh & Liang, 2017), minimax duality and Karush-Kuhn-Tucker (KKT) conditions. Fowl et al. (2021a) used gradient matching which assumes inner optimality has been achieved by a fixed model. Fowl et al. (2021b) avoided the bi-level optimization with a well-trained proxy model from clean data. Different from these previous studies, we propose a gradient-free optimization process to solve the outer optimization pipeline, and show a good optimization result on the attacker’s goal.
@@ Appendix B More Backgrounds of 3D Gaussian Splatting
where $\lambda$ is a trade-off parameter. To ensure finer reconstruction, Gaussian Splatting deploys adaptive density control to automatically add or remove Gaussians from optimization. The process of adding Gaussians, known as densification, involves two primary operations: splitting and cloning. Conversely, removing Gaussians, called pruning, indicates eliminating a Gaussian from the set of optimization variables. Both densification and pruning are non-differentiable operations. The adaptive density control is performed during regular optimization intervals and follows specific, predefined rules. In the initial design (Kerbl et al., 2023), for a Gaussian $G_{i}$ with view-space projection point $\mu^{k}_{i}=(\mu^{k}_{i,x},\mu^{k}_{i,y})$ under viewpoint $V_{k}$, the average view-space positional gradient $\nabla_{\mu_{i}}\mathcal{L}$ is calculated every 100 training iterations:
@@
Appendix C Extended Experiments
C.1 Full attack results
@@
C.2 Full black-box attack results
We report the full results of attacking a black-box victim (Scaffold-GS) (Lu et al., 2024) on NeRF-Synthetic and MIP-NeRF360 datasets in Table 5.
C.3 Ablation study on constraint strength and attack performance
Typically, attacker can choose to balance the attack performance and stealthiness by adjusting attack constraints $\epsilon$. To better understand the tradeoff between $\epsilon$ and attack performance, we poisoned each scene with three levels of $\epsilon$ setting, which are $8/255$, $16/255$ and $24/255$. The comparisons of computation cost under different constraint levels of attacks are provided in Table 6 (NeRF-Synthetic) and Table 7. All results have been reported via the average and standard deviation of three individual runs of victim training.

### v2

@@ Abstract
3D Gaussian splatting (3DGS), known for its groundbreaking performance and efficiency, has become a dominant 3D representation and brought progress to many 3D vision tasks. However, in this work, we reveal a significant security vulnerability that has been largely overlooked in 3DGS: the computation cost of training 3DGS could be maliciously tampered by poisoning the input data. By developing an attack named Poison-splat, we reveal a novel attack surface where the adversary can poison the input images to drastically increase the computation memory and time needed for 3DGS training, pushing the algorithm towards its worst computation complexity. In extreme cases, the attack can even consume all allocable memory, leading to a Denial-of-Service (DoS) that disrupts servers, resulting in practical damages to real-world 3DGS service vendors. Such a computation cost attack is achieved by addressing a bi-level optimization problem through three tailored strategies: attack objective approximation, proxy model rendering, and optional constrained optimization. These strategies not only ensure the effectiveness of our attack but also make it difficult to defend with simple defensive measures. We hope the revelation of this novel attack surface can spark attention to this crucial yet overlooked vulnerability of 3DGS systems. [^1] Our code is available at https://github.com/jiahaolu97/poison-splat.
@@ 1 Introduction
We develop a computation cost attack method Poison-splat as a proof-of-concept of this novel attack vector. Poison-splat takes the form of training data poisoning (Tian et al., 2022), where the attacker manipulates the input data fed into the victim 3DGS system. This is practical in real-world scenarios, as commercial 3D service providers like Kiri (KIRI, ), Polycam (Polycam, ) and Spline (Spline, ) receive images or videos from end users to generate 3D captures. An attacker can hide themselves among normal users to submit poisoned data and launch the attack stealthily, or even secretly manipulate data uploaded by other users. During peak usage periods, this attack competes for computing resources with legitimate users, slowing down service response times and potentially causing severe consequences like service crashes, leading to significant financial losses for service providers.
@@ 4 Experiments
Implementation details. In our experiments, we use the official implementation of 3D Gaussian Splatting [^5] https://github.com/graphdeco-inria/gaussian-splatting (Kerbl et al., 2023) to implement the victim behavior. Following their original implementation, we use the recommended default hyper-parameters, which were proved effective across a broad spectrum of scenes from various datasets. For testing black-box attack effectiveness, we use the official implementation of Scaffold-GS [^6] https://github.com/city-super/Scaffold-GS (Lu et al., 2024b) as a black box victim. All experiments reported in the paper were carried out on a single NVIDIA A800-SXM4-80G GPU.
@@ 4.1 Main Results
We report the comparisons of the number of Gaussians, peak GPU memory, training time and rendering speed between clean and poisoned data across three different datasets in Tables 1 to show the effectiveness of our attack. The full results containing all scenes are listed in the Appendix Table 3 and Table 4. We provide visualizations of example poisoned datasets and corresponding victim reconstructions in Appendix Section F.
@@ 4.2 Black-box Attack Performance
The answer is affirmative. To test the effectiveness of our attack on a black-box victim, we chose Scaffold-GS (Lu et al., 2024b), a 3DGS variant algorithm designed with a focus on efficiency. Scaffold-GS differs significantly from traditional Gaussian Splatting as it utilizes anchor points to distribute local 3D Gaussians, which greatly alters the underlying feature representation. We directly tested Scaffold-GS on the poisoned dataset we crafted for the traditional 3DGS algorithm, making the victim (Scaffold-GS trainer) a black-box model. The results, as presented in Table 2 and Table 5, demonstrate the attack’s capability to generalize effectively, as evidenced by increased Gaussian parameters, higher GPU memory consumption, and reduced training speed.
@@ 5 Conclusion and Discussions
Ethics Statement. While our methods could potentially be misused by malicious actors to disrupt 3DGS service providers and cause financial harm, our aim is not to facilitate such actions. Instead, our objective is to highlight significant security vulnerabilities within 3DGS systems and prompt the community—researchers, practitioners, and providers—to recognize and address these gaps. We aim to inspire the development of more robust algorithms and defensive strategies, enhancing the security and reliability of 3DGS systems in practical applications. We are committed to ethical research and do not endorse using our findings to inflict societal harm.
@@ 6 Acknowledgement
This project is supported in part by NUS Start-up Grant A-0010106-00-00, and by the National Research Foundation, Singapore, and Cyber Security Agency of Singapore under its National Cybersecurity R&D Programme and CyberSG R&D Cyber Research Programme Office (Award: CRPO-GC1-NTU-002).
@@
The Appendix is organized as follows: Section A discusses related works; Section B provides a formal and comprehensive description of 3D Gaussian Splatting training. Section C explicitly describes the threat model in this paper, aiming to provide readers with a clearer understanding. Section D offers the full results of experiments mentioned in the main paper, presented here due to page constraints. Section E lists all the notations used throughout the paper. Section F presents examples of attacker-crafted poisoned images and the corresponding reconstructions from victim models trained with those images. Section G reports the attack performance under various levels of perturbation intensity, ranging from $\epsilon=8/255$ to $\epsilon=24/255$. Section H studies how effective the Poison-Splat attack will be if the poison ratio is not 100%. Section I shows the defensive threshold of limiting Gaussian numbers and discuss the difficulty of setting a uniform defense threshold. Section J discusses another potential defense: image smoothing as a pre-processing, and shows that it is not an ideal defense in terms of keeping reconstruction quality. Section K reports and analyzes the time complexity of Poison-Splat attack.
@@ Appendix A Related Works
3D Gaussian Splatting is receiving overwhelming popularity in the 3D vision domain (Chen & Wang, 2024; Fei et al., 2024; Wu et al., 2024b; Dalal et al., 2024). The explicit nature of 3D Gaussian Splatting enables real-time rendering capabilities and unparalleled levels of control and editability, demonstrating its effectiveness in 3D generation. Since its introduction in (Kerbl et al., 2023), 3D Gaussian Splatting has facilitated a diverse array of downstream applications, including text-to-3D generation (Chen et al., 2024; Tang et al., 2024; Yi et al., 2024), Simultaneous Localization and Mapping (Keetha et al., 2024; Matsuki et al., 2024; Yan et al., 2024a), human avatars (Kocabas et al., 2024; Lei et al., 2024; Zielonka et al., 2023), 4D-dynamic scene modeling (Luiten et al., 2024; Wu et al., 2024a; Yang et al., 2024), and other applications (Xie et al., 2024b; Shen et al., 2024; Ye et al., 2023; Wang et al., 2024b; a).
@@ A.2 Attacks on Machine Learning Systems
Machine learning has been shown to be vulnerable to various types of attacks (Pitropakis et al., 2019; Rigaki & Garcia, 2023; Chakraborty et al., 2018; Oliynyk et al., 2023; Tian et al., 2022; Zhang et al., 2021; 2023). The majority of attacks target the either confidentiality (including membership inference (Shokri et al., 2017; Salem et al., 2019; Choquette-Choo et al., 2021), data reconstruction attacks (Fredrikson et al., 2015; Carlini et al., 2021; 2023; Yu et al., 2023; Geiping et al., 2020; Lu et al., 2022) and model stealing attacks (Tramèr et al., 2016; Chandrasekaran et al., 2020)) or integrity (like adversarial attacks (Biggio et al., 2013; Szegedy et al., 2013; Dong et al., 2018; 2019; Lu et al., 2024a; Zhou et al., ) and data poisoning attacks (Barreno et al., 2010; Jagielski et al., 2018; Biggio et al., 2012; Mei & Zhu, 2015b)). Different from the above mentioned attacks, we aim to illustrate that by carefully crafting input images, the attacker can indeed attack availability, i.e., timely and cost-affordable access to machine learning service. The most similar research is the work by Shumailov et al. (2021), which is the first Denial-of-Service (DoS) attack towards machine learning systems. It proposed sponge examples which can cause much more latency and energy consumption for language models and image classifiers, driving deep learning models to their worst performance. Different from the work (Shumailov et al., 2021), our attack targets 3D reconstruction algorithms and is not driven by neural networks.
A.2.1 Data Poisoning Attacks
@@
We consider our proposed attacks to fall into the category of data poisoning attack, which manipulates training data to launch the attack at the training stage. The attacker makes the first move to craft poisoned training samples, and the victim has a goal of training the system based on samples provided by the attacker. Solving this problem typically involves solving a max-min bi-level optimization problem, where the attacker wants to maximize some adversarial loss (i.e., test error or computational cost) and victim wants to minimize the training objective (i.e., training loss). Different literature solves the problem using different techniques (Biggio et al., 2012; Mei & Zhu, 2015b; a; Li et al., 2016). For many machine learning systems, given the inherent complexity and their reliance on large datasets, it is intractable to exactly do the bi-level optimization. The work (Muñoz-González et al., 2017) exploited back-gradient optimization technique to approximate real gradients through reverse-mode differentiation. The work (Huang et al., 2020) formulated as a meta-learning problem, and approximated the inner optimal solution by training only $K$ SGD steps for each outer objective evaluation. The work (Koh et al., 2022) proposed three methods to approximate the expensive bi-level optimization: influence functions (Koh & Liang, 2017), minimax duality and Karush-Kuhn-Tucker (KKT) conditions. The work (Fowl et al., 2021a) used gradient matching which assumes inner optimality has been achieved by a fixed model. The work (Fowl et al., 2021b) avoided the bi-level optimization with a well-trained proxy model from clean data. Different from these previous studies, we propose a gradient-free optimization process to solve the outer optimization pipeline, and show a good optimization result on the attacker’s goal.
@@ Appendix B More Backgrounds of 3D Gaussian Splatting
where $\lambda$ is a trade-off parameter. To ensure finer reconstruction, Gaussian Splatting deploys adaptive density control to automatically add or remove Gaussians from optimization. The process of adding Gaussians, known as densification, involves two primary operations: splitting and cloning. Conversely, removing Gaussians, called pruning, indicates eliminating a Gaussian from the set of optimization variables. Both densification and pruning are non-differentiable operations. The adaptive density control is performed during regular optimization intervals and follows specific, predefined rules. In the initial design (Kerbl et al., 2023), for a Gaussian $G_{i}$ with view-space projection point $\mu^{k}_{i}=(\mu^{k}_{i,x},\mu^{k}_{i,y})$ under viewpoint $V_{k}$, the average view-space positional gradient $\nabla_{\mu_{i}}\mathcal{L}$ is calculated every 100 training iterations (Ye et al., 2024):
@@
Appendix C Threat Model
We provide a structured description of the attack’s goal, input, output, information and constraints in this section, to provide more clarity of the threat model:
Attacker Goal: To increase the computation cost and over-consume computation resource of 3DGS service provider.
Attacker Input: Clean data.
Attacker Output: Poisoned data by running attack on the clean data.
Attacker Information: Attacker doesn’t need to know the underlying hardware device of the victim.
  • White-box: Attacker knows the specific 3DGS configuration of the victim.
  • Black-box: Attacker only knows victim is using 3DGS, but does not know which variant of 3DGS representation and what configuration the victim is using.
Attacker Constraint: Attacker can optionally choose to constrain the perturbation range $\epsilon$ of his additive perturbation to add on clean data.
Attacker Algorithm: Algorithm 1.
Appendix D Extended Experiments
D.1 Full attack results
@@
D.2 Full black-box attack results
We report the full results of attacking a black-box victim (Scaffold-GS) (Lu et al., 2024b) on NeRF-Synthetic and MIP-NeRF360 datasets in Table 5.
To further test the effectiveness of Poison-Splat, we conduct black-box attack experiments on Mip-Splatting (Yu et al., 2024), which received the best student paper award at CVPR 2024 and has been highly popular (with 178 citations as of November 20th, 2024). Mip-Splatting is an advanced variant of the original 3D Gaussian Splatting, incorporating a 3D smoothing filter and a 2D Mip filter. These enhancements help eliminate various artifacts and achieve alias-free renderings. Results are shown in Table 6.
We found that Mip-Splatting (Yu et al., 2024) consumes more GPU memory compared to the original 3D Gaussian Splatting, making it more prone to the worst attack consequence of running Out-of-Memory. As illustrated in Table 6, even when the attack perturbation is constrained to $\epsilon=16/255$, the GPU memory consumption nearly reached the 80 GB capacity of Nvidia A800. When we apply an unconstrained attack, all scenes in the MIP-NeRF360 dataset will result in denial-of-service.
We examined the code implementation of Mip-Splatting [^7] https://github.com/autonomousvision/mip-splatting, and identified that it uses quantile computation in its Gaussian model densification function  [^8] In Line 524 of mip-splatting/scene/gaussian_model.py which requires massive GPU memory and easily triggers out-of-memory. Our attack highlights the vulnerability in various 3DGS algorithm implementations.